solve the following equation for x
tan^-1(2x) +tan^-1(3x) = n(pi) + 3pi/4

Grade:12

1 Answers

Rinkoo Gupta
9 years ago
tan-1(2x) +tan-1(2y)= n(pi) +3pi/4
tan-1{(2x+3x)/(1-2x.3x)}=n(pi)+3pi/4
5x/(1-6x2)=tan(npi+3pi/4)
5x/(1-6x2)=tan3pi/4=tan(pi-pi/4)=-tanpi/4=-1
5x=-1+6x2
6x2-5x-1=0
x=1,-1/6
Thanks & Regards
Rinkoo Gupta
askIITians Faculty